3. The market for DVDs has supply and demand curves given by P=2Q and P=42-4 respectively.
Answer #1

Answer

P=$20

quantity demanded is

20=42-Q

Q=42-20=22

quantity supplied is

20=2Q

2Q=20

Q=10

the quantity demanded is higher than the quantity supplied so there is a shortage in the market

shortage =Qd-Qs=22-10=12 units

the shortage is 12 units